OpenStudy (anonymous):

Rewrite the equation of a line in point-slope form to slope intercept form: y-2- 2/3(6x+9)

hartnn (hartnn):

is that y-2 = 2/3(6x+9) ? start by adding 2 on both sides, what u get ?

OpenStudy (anonymous):

y+2= 2/3 (6x+9) +2?

hartnn (hartnn):

on left side, y+2-2 , so only 'y' remains, right ? y = 2/3 (6x+9) +2 ok ?

hartnn (hartnn):

now distribute the 2/3 to both terms inside bracket

OpenStudy (anonymous):

wait how do i distribute it? ._.

hartnn (hartnn):